### v3

# Auto-ARGUE: LLM-Based Report Generation Evaluation

Abstract
Generation of long-form, citation-backed reports is a primary use case for retrieval augmented generation (RAG) systems. While open-source evaluation tools exist for various RAG tasks, ones tailored to report generation are lacking. Accordingly, we introduce Auto-ARGUE, a robust LLM-based implementation of the recent ARGUE framework for report generation evaluation. We present analysis of Auto-ARGUE on the report generation pilot task from the TREC 2024 NeuCLIR track, showing good system-level correlations with human judgments. We further release a web app for visualization of Auto-ARGUE outputs.
1 Introduction
As RAG applications have proliferated, interest in automatic evaluation methodologies for RAG tasks has grown in tandem. Many works have positioned themselves as general solutions for RAG evaluation [2, 3, 8, 10], and while some evaluation desiderata are shared across tasks, task-dependent considerations persist.
Here, we focus on report generation (RG), a RAG task that aims to produce a long-form, citation-attributed response to a complex user query. Two key features distinguish RG from related RAG tasks—notably, long-form QA. First, RG strongly foregrounds the identity of the user (or requester): the same query should yield different reports for requesters with different levels of education or domain expertise. Second, the ideal report represents a summary over the entire corpus of the most user-critical information, thus stressing coverage where traditional QA emphasizes mere adequacy of the response.
While the “task-agnostic” RAG evaluation methodologies discussed above could be applied to RG, our work focuses on the ARGUE framework [7]—the only framework designed expressly for RG—and presents three contributions:
  1. 1.
    Auto-ARGUE: An automatic, LLM-based Python implementation of the ARGUE framework—the only publicly available ARGUE implementation.
  2. 2.
    ARGUE-viz: A web app for visualizing Auto-ARGUE outputs.
  3. 3.
    Case Study: Meta-evaluation results with Auto-ARGUE on the TREC 2024 NeuCLIR report generation pilot task [5].
Auto-ARGUE is configurable, easy-to-use, and extensible to other RG datasets, adopting the new TREC format for RAG outputs [^1] Format schema and validator: https://github.com/hltcoe/rag-run-validator. We release Auto-ARGUE  [^2] https://github.com/hltcoe/auto-argue and ARGUE-viz  [^3] https://github.com/hltcoe/argue-viz to facilitate further work on automatic RG evaluation.
2 System Overview
2.1 Framework: ARGUE
Figure 1: The ARGUE framework from [7], adapted with permission.
Figure 1 depicts the ARGUE framework, which evaluates reports via a tree of binary, sentence-level judgments (blue diamonds) about each sentence’s content and citations. Depending on the path(s) traversed for each sentence, a report may incur penalties (red circles), rewards (green), or neither (beige).
Inputs. ARGUE takes as input: (1) the generated report; (2) the report request, including a problem statement describing the information need and a user story describing the requester; (3) the document collection used to generate the report (with optional relevance judgments); and (4) a collection of nuggets (QA pairs), with links from nugget answers to documents that attest them.
Content Evaluation. Following much prior work [11, 6, 9, 8, 1], ARGUE evaluates reports’ coverage of relevant information via sets of nuggets—QA pairs that represent key questions an ideal report would address, paired with answers linked to documents in the collection that attest them. Key questions that are unanswerable from the collection can also be represented as nuggets with empty answer sets. Each report sentence is assessed to determine which nugget question(s) it answers correctly and reports are rewarded for each such nugget.
Citation Evaluation. Citations are assumed to support only the sentence they are attached to. Sentences may bear $\geq 0$ citations. Relevant citations that attest a sentence are rewarded; non-attesting or missing citations are penalized.
Metrics. ARGUE is agnostic to the magnitudes of its rewards and penalties and to the metrics to be reported based on them. However, Mayfield et al. recommend two metrics—sentence precision and nugget recall—discussed below.
We refer readers to [7] for further details on ARGUE.
2.2 Implementation: Auto-ARGUE
ARGUE leaves much to implementation, such as the judge, the magnitudes of rewards/penalties, the source of nuggets and relevance assessments, and the metrics to report. Here, we detail the choices we made with Auto-ARGUE.
  • LLM Judge. An LLM judge is queried via few-shot prompts to obtain binary (YES/NO) answers to all non-trivial judgments (starred in Figure 1) for a report sentence. Answers to other judgments are determined via lookup.
  • Relevance. Auto-ARGUE deems a document relevant iff it attests some nugget answer, determined via lookup in the nugget set or via the LLM.
  • Nuggets. Nuggets may have multiple answers (each of which may be attested by $\geq 1$ document(s)) and come in two varieties: “AND” nuggets, for which all answers must be given, and “OR” nuggets, for which only one answer (of several) is required. Nuggets may also have importance labels (“vital” or “okay”). Answer attestation is assessed per-sentence but answers are aggregated across all sentences to identify correctly answered nuggets.
  • Metrics. Auto-ARGUE implements the two metrics suggested by Mayfield et al. [7]. Sentence precision is the proportion of sentences that are attested by each of their citations. Nugget recall is the proportion of nuggets correctly answered by the report, with a weighted variant that weights nuggets by importance (okay=1.0; vital=2.0). Auto-ARGUE also computes (un)weighted F1 score(s) based on these two metrics, which can serve as an overall score for a report. Auto-ARGUE also outputs several other fine-grained metrics.
2.3 Visualization: ARGUE-viz
ARGUE-viz is a simple Streamlit [^4] https://streamlit.io app for visualizing Auto-ARGUE outputs for individual runs, including both metrics and fine-grained judgments. Users can toggle between per-topic results and aggregated metrics via radio buttons on a sidebar. Per-topic results display core (e.g. sentence precision, nugget recall) and non-core metrics and statistics (e.g. % relevant citations) for that topic as well as detailed judgment information about the report generated for that topic. Judgment information is displayed via two views: a report view that shows report-level information about supported sentences and correctly answered nuggets and a sentence view that shows similar information at the sentence level (e.g. which nugget answers are (not) attested by that sentence). Collectively, these features enable fine-grained human analysis of errors to facilitate system development.
3 Case Study: TREC NeuCLIR 2024 Report Generation
We evaluate Auto-ARGUE on the 51 runs from the RG pilot task of the TREC 2024 NeuCLIR track [5], which requires generating English reports from one of three non-English collections (Chinese, Russian, Farsi—17 runs each). Human assessors judged sentence support and nugget recall on reports for the same 21 topics for each run. Each topic has 10-20 nuggets, and assessors also identified documents attesting each answer, thus providing the set of relevant documents.
Separately, we obtain the same metrics from Auto-ARGUE, using Llama-3.3 70B as the LLM judge [4] for D, C, G, and H judgments in Figure 1. B judgments use the human relevance assessments. Since all NeuCLIR nuggets are answerable, E and F judgments are not generated. We obtain system rankings from (a) assessor-based and (b) Auto-ARGUE-based macro-average sentence precision and nugget recall across all topics for each language. We compute agreement between these rankings using (i) Kendall’s tau and (ii) accuracy w.r.t. whether two Wilcoxon tests (one per ranking) for each pair of systems $s_{1}$ and $s_{2}$ agree. Figure 2 presents the results. Broadly, we observe good agreement between the two rankings on both metrics, with particularly strong results on sentence precision. More capable LLM judges could yield still further improvements.
Figure 2: Auto-ARGUE vs. human agreement on system rankings based on sentence precision (left) and nugget recall (right) for the TREC 2024 NeuCLIR RG pilot task.
4 Conclusion
This work has introduced Auto-ARGUE—a robust, configurable, LLM-based implementation of the ARGUE framework for report generation (RG) evaluation—as well as ARGUE-viz—a simple Streamlit application for visualization of Auto-ARGUE outputs. Analysis of Auto-ARGUE on the TREC 2024 NeuCLIR RG pilot task with a relatively small open-source LLM judge (Llama-3.3 70B) reveals good correlations with human judgments on system rankings based on the key metrics of sentence precision and nugget recall. We release both Auto-ARGUE and ARGUE-viz to facilitate future work on RG evaluation.
{credits}
4.0.1 Acknowledgements
The authors thank all the participants of the SCALE 2025 workshop at the JHU HLTCOE for valuable feedback on, and testing of, Auto-ARGUE.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.